Oh boy! I just don't get it.
1/n+ 1/(n+1)= -1/n(n+1)

Answer by Alan3354(21584)

1/n+ 1/(n+1)= -1/n(n+1)
Add the 2 fractions on the left. To do that, get a common denominator
DEN = n(n+1), the same on the right side.
---------
(n+1)/DEN + n/DEN = -1/DEN
Multiply thru by DEN
(n+1) + n = -1
2n+1 = -1
2n = -2
n = -1